OpenStudy (anonymous):

Can someone help me with understanding the change of base formula? I am supposed to find the logarithm using the change of base formula for this question log5to the power of 80?

OpenStudy (anonymous):

Log (base 5) 80 ??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I guess I know how to switch the numbers over to replace in the formula but I keep getting large numbers when I put it in the calculator.

OpenStudy (anonymous):

In change of base, You can Select any arbitrary base x (The Selection of x should be such that it helps in simplification) and above Log(base5) 80 = Log (base5) 16*5 = Log (base5) 16 + Log (base5) 5 = Log (base5) 16 +1 = (Log (base2) 16)/(Log (base2) 5) +1

OpenStudy (anonymous):

=(Log (base2) 2^4)/(Log (base2) 5) +1 = (4Log (base2) 2)/(Log (base2) 5) +1 = (4)/(Log (base2) 5) +1 Now I guess i am stuck at (Log (base2) 5). so it all depends on the base to be chosen as in i chosed 2 here...

OpenStudy (anonymous):

I am using the base of 10 would you be able to help me calculate log (base 10) 80 divided by log(base10) 5

OpenStudy (anonymous):

Yes, You can separate the terms as in lets mark x= log (base 10) 80 and y =log(base10) 5 We solve the x first. 10^x =80 for y . 10^y =5

OpenStudy (anonymous):

are you still there cuz I disconnected somehow but I see your answer. This is my first time trying this sight.

OpenStudy (anonymous):

yeah I am here.Hmm Let me think a way out to solve this.

OpenStudy (asnaseer):

if:\[\log_ab=c\]then this implies:\[b=a^c\]do you understand that rule?

OpenStudy (anonymous):

Hey I figured it out. Instead of trying to put it in the calculator as log base and # I was supposed to put it as ln(80)/ln(5) and I got the answer when I did that.

OpenStudy (asnaseer):

if you do, then you can use that result to derive the log to any other base.

OpenStudy (anonymous):

yeah I do understand that part thank you for both of your help.

OpenStudy (anonymous):

Hahaa...Ohh Yes If You Could Use the Calculator You should have told me..lol. I was trying to solve it without it. And You gt the answer, because in Calculator every log is base to 10 or e(exponential)

OpenStudy (anonymous):

Ive never done this before am I supposed to click best answer or something.

OpenStudy (asnaseer):

so taking the last result, if we take log to the base d of both sides (where d is the new base we want), then we get:\[b=a^c\]\[\log_db=\log_d(a^c)=c*\log_da\]therefore:\[c=\frac{\log_db}{\log_da}\]therefore:\[\log_ab=\frac{\log_db}{\log_da}\]

OpenStudy (asnaseer):

this is known as the change of base formula

OpenStudy (asnaseer):

I'm not here for medals - more for teaching - so feel free to award medal to whomever you feel like :)